A professor of History is teaching a section of 100 students. Her first exam’s grade distribution follows.
Exam grades

Frequency

45 to < 50

1

50 to < 55

2

55 to < 60

6

60 to < 65

19

65 to < 70

12

70 to < 75

22

75 to < 80

12

80 to < 85

13

85 to < 90

11

90 to < 95

0

95 to < 100

2

Calculate the standard deviation for this grouped data.
1
Expert's answer
2020-08-25T16:05:53-0400

The formula for standard deviation ("\sigma") for grouped data is given by,


"\sigma = \sqrt{\frac{\sum f*(m-\bar x)^2}{\sum f}}"


where,

f = frequency of the class

m = class midpoint

"\bar x" = mean


and the formula for the mean "\bar x = \frac{\sum(f*m)}{\sum f}"


From the given data we can find the values of above variables as under-





Hence we get,


mean "\bar x = 72"


standard deviation "\sigma = 10.1488"